OpenStudy (summersnow8):

Use the following pKa values to answer questions the next four questions. Lactic acid pKa = 3.86, TrisH+ (“Tris acid”) pKa = 8.10, NH4+ pKa = 9.25, Formic acid pKa = 3.75, H3 PO4 pKa = 2.12, H2 PO4- pKa = 6.82, HPO42- pKa = 12.4 What is the ratio of the conjugate base to conjugate acid in the following buffer solutions? 1) Formate buffer, pH = 4.00 2) 1.0 M Phosphate buffer, pH = 7.00 3) 1.0 M Tris buffer, pH = 7.00 4) 0.10 M Tris buffer, pH = 7.00 5) Ammonia/Ammonium

OpenStudy (summersnow8):

I am unsure about 2 - 4...... My work for 1) & 5): 1) Formate buffer, pH = 4.00 pH = pKa + log [base]/[acid] 0.25 = log R pH = pKa + log R 100.25 = R 4.00 = 3.75 + log R 1.78 / 1 = R 5) Ammonia/Ammonium buffer, pH = 9.60 pH = pKa + log [base]/[acid] 0.25 = log R pH = pKa + log R 101.5 = R 9.60 = 8.10 + log R 31.6 / 1 = R

OpenStudy (caozeyuan):

for 2, whats your conjugate acid, and conjugate base(hint: compare pKa of buffer to the pKa of your CA)

OpenStudy (caozeyuan):

also, know that 1,0M is your total concentration

OpenStudy (summersnow8):

so I could use H3PO4 for my acid and H2PO4 for my base?

OpenStudy (summersnow8):

I am not really sure how I factor in concentration

OpenStudy (summersnow8):

Is my work correct for 1 & 5?